a container is full of oil and has a square base of side 0.6m and a height of 0.275m.cylinders have a radius of 6cm and height of 18 these are filled from the container. calculate the volume of the container and cylinder and the number of cylinders which can be filled
Answer by Theo(13342)

Before you start, you should convert all measurements to the same unit.

the container is in meters and the cylinders are in centimeters.

I would convert all measurements to centimeters to make them consistent.

Since 1 meter = 100 centimeters, the measurements of the container become:

square base side of .6m * 100 = 60cm
height of .275m * 100 = 27.5cm

The volume of the container is height * area of the base.

Area of the base = 60*60 = 3600 square centimeters.

Volume of the container = 3600 * 27.5 = 99000 cubic centimeters.

The cylinders have a radius of 6 centimeters and a height of 18 centimeters.

The area of the base = pi*r^2 = pi*6^2 = pi*36 = 113.09733565 square centimeters.

The volume of the cylinder is equal to the area of the base times the height which equals 113.09733565 * 18 = 2035.75204 cubic centimeters.

Since the volume of the container is 99000 cubic centimeters and the volume of each cylinder is 2035.75204 cubic centimeters, then you divide the volume of the container by the volume of each cylinder to get a total of 48.63067706 cylinders that can be filled by the container.

Round that to the nearest lowest integer and you get 48 cylinders can be filled by the container with the 49th cylinder being only partially filled.